# Modeling Short-Range Microwave Networks to Scale Superconducting Quantum Computation

### v3

@@ IV.3 Modeling fragments of computation
Specifically, local CX gate noise is chosen to range from 0.0045 to 0.0395 in increments of 0.025 for the “monolithic” case, extrapolating upward from the hybrid, Surface and Bacon-Shor threshold as considered in IV.2.1. Link noise is chosen to apply a 1% depolarizing channel followed by an amplitude-damping channel with $\eta=0.12$ based on channels modeled in Subsection III.3. The results appear as Figure 10 and show a qualitatively similar pattern to those of Subsection IV.2. The almost perfectly linear appearance of the plot is due to Qiskit’s idealized density matrix comparison, which does not introduce measurement randomness as would tomography.
IV.4 Detecting errors in links
@@
is achievable, implying asymptotically high efficiency upon success with scaling of $m$ in return for a (detected) failure probability that scales with $k$.
Figure 11: Basic error detection for amplitude-damping links. The “$+$” symbols represent $X$ and $CX$ gates, the “$I$” a trivial identity gate, and the two maroon boxes two copies of the link.
@@
a qubit version of the [111, Equation 2.15]. This is illustrated via Figure 11 A slightly more sophisticated code on 2 qubits is given by
@@
Qubits Sent Qubits Used Error Multiplier Efficiency
@@
Table 2: Table of encodings for up to 10 qubits. The error multiplier is the number of “1”s in the computational basis, such that this number times $\eta$ is the total probability of a detected failure. The listed efficiency is the number of qubits sent divided by the number used, conditioned on successful transmission.
@@
With a CX error of about 0.01 depolarizing noise and a 0.02 readout error, we may roughly model the conditional channel applied on a “success” measurement as probably about 0.03-0.04 depolarizing with current local error rates. The 0.88 reported efficiency from [52] and 0.02 readout error yield a post-selection efficiency of about 0.86. As mentioned in that paper, however, it should be possible to increase the efficiency in future experiments. Similarly, [53] suggests a potential transfer fidelity of 0.96. Error detection with post-selection is potentially most useful in relatively small, early distributed experiments. Post-selection is not ideal for long SWAP chains with many potential failure points, but it may perform better in the tree or expander connectivities as proposed in Section IV.1. Logarithmic depth circuits with expanderized connectivity woud have inverse polynomial success probability even without retrying any transfers mid-computation.
Broadly, we expect on-chip noise to fall to extremely small values with continued engineering research. If the fidelity of link transfers plateaus at a lower value, such as the 0.96 mentioned in  [53], then comparisons such as those in Section IV would ultimately show a disadvantage to using links with extremely high fidelity of local processing. In this regime, error detection may enable continued advantages in chiplet architectures. With sufficiently good on-chip processing and readout, one may consider an abstracted transfer operation that distributes entanglement, detects errors until retrial, and ultimately teleports the computation qubits. Such a scheme may still have a smaller on-chip footprint than error correction, requiring only two qubits in principle. Hence link error detection could be most valuable in a regime of small, high-quality chiplets.
V Discussion and Conclusions
@@
    In Section IV, we compare possibilities for state-of-the-art microwave links to connect current, industrial-quality chips. A major experimental question is how and if these technologies will be combined. To account for possible impacts of integration and unpredictability of future developments, we consider a wide range of possible fidelities for both on-chip operations and links. An essential next step in distributed quantum computing is experimental work to begin resolving these ambiguities.
@@
    The analysis of Section IV.2 hints that for plausible ranges of noise parameters, the overall error probability for simultaneous operations linking many qubits will be lower in a chiplet than in a monolithic architecture. Intuitively, a logical qubit spread over a chiplet system could have looser fault-tolerance requirements than one implemented monolithically. Quantum error correction is a major topic of its own with many complexities, options, and subtleties, so it is beyond the scope of this paper to rigorously investigate this intuition. Nonetheless, it seems likely that in contrast to common expectations, distribution may usefully lie below the logical qubit layer in a fault-tolerant scheme.
@@
    Today’s optical and microwave links represent vastly different regimes in expected range and performance as described in Section III. There is nonetheless optimism that in the long term, matter-optical transduction will benefit from major advances in hardware and local error reduction. It remains open if local networks could eventually use optical fiber to similar effects as proposed for microwave links in this work.

### v4

@@ 4.3 Modeling fragments of computation
Specifically, local CX gate noise is chosen to range from 0.0045 to 0.0395 in increments of 0.025 for the “monolithic” case, extrapolating upward from the hybrid, Surface and Bacon-Shor threshold as considered in 4.2.1. Link noise is chosen to apply a 1% depolarizing channel followed by an amplitude-damping channel with $\eta=0.12$ based on channels modeled in Subsection 3.3. The results appear as Figure 10 and show a qualitatively similar pattern to those of Subsection 4.2. The almost perfectly linear appearance of the plot is due to Qiskit’s idealized density matrix comparison, which does not introduce measurement randomness as would tomography.
4.4 An error detection scheme for links
@@
is achievable, implying asymptotically high efficiency upon success in return for a (detected) failure probability that scales with $k$.
Figure 11: Basic error detection for amplitude-damping links. The “$+$” symbols represent $X$ and $CX$ gates, the “$I$” a trivial identity gate, and the two dashed boxes are two copies of the link.
@@
a qubit version of the [106, Equation 2.15]. This is illustrated via Figure 11 A slightly more sophisticated code on 2 qubits is given by
@@
Q Sent Q Used Err Mult Efficiency
@@
Table 2: Table of encodings for up to 10 qubits. “Q Sent” and “Q Used” refer respectively to the number of qubits sent and the number used (including those for error detection). The error multiplier is the number of “1”s in the computational basis, such that this number times $\eta$ is the total probability of a detected failure. The listed efficiency is the number of qubits sent divided by the number used, conditioned on successful transmission.
@@
With a CX error of about 0.01 depolarizing noise and a 0.02 readout error, we may roughly model the conditional channel applied on a “success” measurement as probably about 0.03-0.04 depolarizing with current local error rates. The 0.88 reported efficiency from [52] and 0.02 readout error yield a post-selection efficiency of about 0.86. As mentioned in that paper, however, it should be possible to increase the efficiency in future experiments. Similarly, [53] suggests a potential transfer fidelity of 0.96. Error detection with post-selection is potentially most useful in relatively small, early distributed experiments. Post-selection is not ideal for long SWAP chains with many potential failure points, but it may perform better in the tree or expander connectivities as proposed in Section 4.1. Logarithmic depth circuits with expanderized connectivity woud have inverse polynomial success probability even without retrying any transfers mid-computation.
Broadly, we expect on-chip noise to fall to extremely small values with continued engineering research. If the fidelity of link transfers plateaus at a lower value, such as the 0.96 mentioned in  [53], then comparisons such as those in Section 4 would ultimately show a disadvantage to using links with extremely high fidelity of local processing. In this regime, error detection may enable continued advantages in chiplet architectures. With sufficiently good on-chip processing and readout, one may consider an abstracted transfer operation that distributes entanglement, detects errors until retrial, and ultimately teleports the computation qubits. Such a scheme may still have a smaller on-chip footprint than error correction, requiring only two qubits in principle. Hence link error detection could be most valuable in a regime of small, high-quality chiplets.
5 Discussion and Conclusions
@@
    In Section 4, we compare possibilities for state-of-the-art microwave links to connect current, industrial-quality chips. A major experimental question is how and if these technologies will be combined. To account for possible impacts of integration and unpredictability of future developments, we consider a wide range of possible fidelities for both on-chip operations and links. An essential next step in distributed quantum computing is experimental work to begin resolving these ambiguities.
@@
    The analysis of Section 4.2 hints that for plausible ranges of noise parameters, the overall error probability for simultaneous operations linking many qubits will be lower in a chiplet than in a monolithic architecture. Intuitively, a logical qubit spread over a chiplet system could have looser fault-tolerance requirements than one implemented monolithically. Quantum error correction is a major topic of its own with many complexities, options, and subtleties, so it is beyond the scope of this paper to rigorously investigate this intuition. Nonetheless, it seems likely that in contrast to common expectations, distribution may usefully lie below the logical qubit layer in a fault-tolerant scheme.
@@
    Today’s optical and microwave links represent vastly different regimes in expected range and performance as described in Section 3. There is nonetheless optimism that in the long term, matter-optical transduction will benefit from major advances in hardware and local error reduction. It remains open if local networks could eventually use optical fiber to similar effects as proposed for microwave links in this work.